A recent SBPA grad has accepted a job as a DC Realtor. She eatimated her gross pay for the next 3 years is 45,000 in year 1, 56,000 in year 2 , and 61,000 in year 3. What is the present value of these cash flows if they are discounted at 3/%?

Solomon intends to deposit $2500 per year at the end of the year for the next 20 years into a Roth IRA that pays 6% he plans to withdraw the tax- free balance at the end of the 20th year. How much would that balance be?

Suppose you have a winjing DC lottery ticket which pays the following sums at the end of each of the next ten years : 16,000 a year for the first four years, $18,000 each year for the next four years , and 25,000 in years nine and ten. what is the present
value of these payouts?

return to question C What is the future value of these payouts (i.e. at the end of the year 10?
Answer #1

Solution:

1.

It is given in the question that the cash flow is as follows 45,000 in Year 1,  56,000 in year 2 , and 61,000 in year 3.

The given discount rate is 3%

Present value can be calculated by discounting the future cash flow with discount rate.

Present Value = 45,000 /(1+0.03)^1 + 56,000/(1+0.03)^2 + 61,000/(1.03)^3 = 43689.32 + 52785.37 +55823.64 = 152,298.33

2.

Payment per period = $2500, Interest rate is 6%, Period = 20

In the given question we need to calculate the future value of annuity

FV = P (1+r) - 1

FV = 2500 * (1 +0.0620 - 1 0.06

2.2071 FV = 2500 * 0.06

FV = 91,963.98

The Balance would be 91,963.98

In question 3 the interest rate is not given.